# On-the-fly Point Feature Representation for Point Clouds Analysis

### v1

@@ 1. Introduction
Point cloud analysis on robotics and automation application (Zermas et al., 2017; Wu et al., 2019b; Wiesmann et al., 2022; Nunes et al., 2022; Jie et al., 2023) has garnered substantial attention in recent years, driven by advancements in sensor technologies like LiDAR and photogrammetry. This growing interest attributes to two key advantages: 1) It can accurately represent complex objects with numbers of points. 2) It can be quickly created by using 3D scanning devices. Compared to 2D image data, point clouds provide a more powerful 3D sparse representation containing abundant geometry and layout information of the environment.
@@
Local geometric structures are vital for understanding point clouds. In an effort to capture this information, some prior works attempt to learn local relationships from convolutions (Li et al., 2018; Wu et al., 2019a; Jiang et al., 2018), attentions (Guo et al., 2021; Zhao et al., 2021; Yu et al., 2022), or graphs (Yang et al., 2018; Wang et al., 2019; Zhang et al., 2019). However, these methods require huge amount of labelled data to learn local geometry implicitly (Ran et al., 2022), while getting large amount of labelled 3D annotations is difficult. Recently, RepSurf (Ran et al., 2022) has emerged as a novel approach that explicitly learns geometric information based on umbrella surface (Foorginejad and Khalili, 2014), which is a triangle set [^1] In this paper, we refer triangle set to a collection of connected or disconnected triangles. For those connected ones, they form a surface. with connected triangles formed by $k$ nearest neighbors ($k$-NNs). While triangle sets are effective in capturing location and orientation information, they often fall short in incorporating curvature knowledge, which is essential for accurate point cloud recognition (Sun et al., 2016; Czerniawski et al., 2016). Moreover, as depicted in the supplementary material, in certain $k$-NNs, the $k$ points may come from different surfaces of the object. These “noisy points” can lead to the distortion of $k$-NN triangle sets, significantly impacting the quality of the obtained geometric features (Ran et al., 2022).
@@
Moreover, our OPFR is backbone-agnostic, making it compatible with different 3D point clouds analysis architectures. We demonstrate its model-agnostic nature by adapting two representative backbones: PointNet++ (Qi et al., 2017b) and Point Transformer (Zhao et al., 2021). It serves as an efficient plug-and-play module, and achieves substantial performance improvements. Empirical results prove its compatibility with different backbones. When incorporating with Point Transformer backbone, our OPFR achieves state-of-the-art performance for both point cloud classification and semantic segmentation tasks.
@@ 2.1. Deep Learning on Point Clouds
As PointNet++ establishes the hierarchical point clouds analysis framework, the focus of many works has shifted towards the development of local feature extractors, including convolution-based (Li et al., 2018; Jiang et al., 2018; Mao et al., 2019; Komarichev et al., 2019; Wu et al., 2019a), attention-based (Guo et al., 2021; Zhao et al., 2021; Yu et al., 2022), and graph-based (Yang et al., 2018; Te et al., 2018; Wang et al., 2019; Zhang et al., 2019; Xu et al., 2020) approaches. PointCNN (Li et al., 2018) learns a $\chi$-transformation from input point clouds, which attempts to re-organize inputs into canonical order. Subsequently, it utilizes vanilla convolution operations to extract local features. Point Transformer (Zhao et al., 2021) replaces conventional shared-MLP modules with Transformer (Vaswani et al., 2017) blocks, serving as feature extractors within localized patch processing. DGCNN (Wang et al., 2019) utilizes dynamic graph structures to enhance feature learning and capture relationships between points. However, these works rely heavily on the learnability of feature extractors, potentially missing inherent local shape information. More recently, RepSurf (Ran et al., 2022) leverages triangle sets with connected triangles formed by $k$ nearest neighbors ($k$-NNs), to learn location and orientation-aware representations from geometric features explicitly. Although location and orientation features are explicitly injected into network architecture in RepSurf, the usage of curvature information still remains underexplored. Moreover, RepSurf relies on naive $k$-NNs to produce triangle sets and obtain geometric features, which are vulnerable to noisy points (Ran et al., 2022).
@@ 3. Methodology
The pipeline of On-the-fly Point Feature Representation (OPFR) is depicted in Fig. 1, and we illustrate the OPFR generating process for the right corner of table (highlighted in pink), which is one of our interested points. Firstly, we propose Hierarchical Sampling module, which takes each point in the point cloud as input and outputs several clusters (highlighted in blue) and corresponding centroids (highlighted in orange). This hierarchical sampling strategy improves the quality of triangle set for each point, thereby facilitating the development of subsequent geometric features. Then, for each point pair within the clusters, we design Curve Feature Generator module to generate geometric features, including location, orientation, and curvature. The inclusion of explicit curvature information allows us to more effectively capture the local geometry surrounding these point pairs. To enhance the efficiency and enable on-the-fly processing, we present Local Reference Constructor module. It approximates a local coordinate system (highlighted in red) for each point pair using adjacent points from triangle sets. Lastly, these obtained geometric features are further fed into shared-MLP followed by a pooling operation, constituting our final OPFR representation $\mathbf{r}_{i}$. The resultant OPFR representation $\mathbf{r}_{i}$ along with coordinate $\mathbf{x}_{i}$ can be directed into various point clouds analysis backbones, e.g., PointNet++ (Qi et al., 2017b) and Point Transformer (Zhao et al., 2021), for end-to-end training.
@@ 3.3. Curve Feature Generator
PointTrans. (Zhao et al., 2021) 1k pnts 93.7 90.6 82.3 80.7 5.187M 0.29G
@@ 3.4. On-the-fly Point Feature Representation
In Fig. 3, the three-view drawing depicts the OPFR values of 1-st channel for an airplane. The blue hues represent areas with smaller OPFR values, typically the airplane body, while the red hues indicate larger OPFR values, primarily associated with the airplane wings. This color differentiation underlines our OPFR is sensitive to curvature variation across the airplane’s structure, demonstrating the curvature-aware property of OPFR. We provide more visualization examples in the supplementary material. Additionally, it is important to highlight that, as shown in Fig. 4, our OPFR can outperform vanilla PFH by a large margin with the help of learnable shared-MLP. Furthermore, the introducing of shared-MLP only increases 0.012M learnable parameters, which is approximately negligible for most popular backbones (Qi et al., 2017b; Zhao et al., 2021).
@@ 4. Experiments
We evaluate our OPFR on two primary tasks: point cloud classification and semantic segmentation. We choose two representative point cloud understanding models, PointNet++ (Qi et al., 2017b) and Point Transformer (Zhao et al., 2021), as our backbones to evaluate the effectiveness and compatibility of OPFR representations across different backbone architectures. Additionally, we carry out ablation studies to demonstrate the effectiveness of our OPFR network designs and quantitatively evaluate the efficiency and quality of OPFR feature representations. Moreover, due to space constraints, we present qualitative results in the supplementary material.
@@ 4.1. Classification
Classification on ModelNet40. ModelNet40 (Wu et al., 2015) is one synthetic object classification benchmark, which contains 9843 training samples and 2468 testing samples. They contain 100 unique CAD models from 40 object categories. The experimental results are presented in Tab. 1. The results reveal that our OPFR significantly improves PointNet++ (Qi et al., 2017b) backbone by 3.8% OA and 3.2% mAcc, with just an additional 0.012M more parameters and 0.15G more FLOPs. The scaled-up OPFR further attains a slight improvement of 0.1% OA and 0.2% mAcc. Moreover, when integrated with transformer-based backbone, Point Transformer (Zhao et al., 2021), our OPFR achieves the state-of-the-art 94.8% OA and 92.0% mAcc (+1.1% OA and +1.4% mAcc).
@@ 4.2. Semantic Segmentation
Semantic Segmentation on S3DIS. S3DIS (Armeni et al., 2016) encompasses 271 scenes which are distributed across 6 indoor areas, with each individual point being classified into one of 13 semantic labels. Following a common protocol (Tchapmi et al., 2017; Qi et al., 2017b), we evaluate the presented approach in two modes: (a) Area-5 is withheld for training and is used for testing, and (b) 6-fold cross-validation. In Tab. 2, our proposed OPFR considerably enhances PointNet++ (Qi et al., 2017b) by 14.7%/16.9%/3.0% (mIoU/mAcc/OA) on S3DIS 6-fold benchmark. Our result is comparable to PointNeXt-XL (Qian et al., 2022), with around 40$\times$ fewer parameters and 11$\times$ fewer FLOPs. When integrated with Point Transformer (Zhao et al., 2021), the performance of OPFR exceeds previous state-of-the-art Superpoint Transformer (Sun et al., 2023) by 0.9%/0.1%/1.6% (mIoU/mAcc/OA) for S3DIS 6-fold. Meanwhile, on S3DIS Area-5, our OPFR attains mIoU/mAcc/OA of 72.6%/78.6%/91.7% (+2.2%/+2.1%/+0.9%), surpassing the prior state-of-the-art ConDaFormer (Duan et al., 2024).
@@ 5. Conclusion
We propose the novel plug-and-play module On-the-fly Point Feature Representation (OPFR) for various backbones. It explicitly captures local geometry including location, orientation and curvature through Curve Feature Generator module. We further develop the Local Reference Constructor module to improve efficiency and enable on-the-fly processing. Additionally, we introduce the Hierarchical Sampling module to mitigate the distortion of triangle sets that occurs in the naive $k$ nearest neighbors sampling, thereby enhancing the robustness of obtained geometric features. We evaluate the proposed OPFR on ModelNet40 (Wu et al., 2015) and ScanObjectNN (Uy et al., 2019) benchmarks for point cloud classification task, S3DIS (Armeni et al., 2016) for semantic segmentation task. For both PointNet++ (Qi et al., 2017b) and Point Transformer (Zhao et al., 2021) backbones, our presented OPFR achieves the state-of-the-art results on different benchmarks. The comprehensive empirical results demonstrate the backbone-agnostic nature of our proposed method. We believe that our work can prompt consideration of how to better leverage geometric knowledge in network architecture designs for understanding point clouds.

### v2

@@ 1. Introduction
Point cloud analysis on robotics and automation application (Chen et al., 2024; Cheng et al., 2023; Zhao et al., 2021a; Sheng et al., 2022; Wiesmann et al., 2022; Zhao and Lee, 2022; Sheng et al., 2023; Han et al., 2024; Li and Zhao, 2024; Li et al., 2024) has garnered substantial attention in recent years, driven by advancements in sensor technologies like LiDAR and photogrammetry. This growing interest attributes to two key advantages: 1) It can accurately represent complex objects with numbers of points. 2) It can be quickly created by using 3D scanning devices. Compared to 2D image data, point clouds provide a more powerful 3D sparse representation containing abundant geometry and layout information of the environment.
@@
Local geometric structures are vital for understanding point clouds. In an effort to capture this information, some prior works attempt to learn local relationships from convolutions (Li et al., 2018; Wu et al., 2019; Jiang et al., 2018), attentions (Guo et al., 2021; Zhao et al., 2021b; Yu et al., 2022), or graphs (Yang et al., 2018; Wang et al., 2019; Zhang et al., 2019). However, these methods require huge amount of labelled data to learn local geometry implicitly (Ran et al., 2022), while getting large amount of labelled 3D annotations is difficult. Recently, RepSurf (Ran et al., 2022) has emerged as a novel approach that explicitly learns geometric information based on umbrella surface (Foorginejad and Khalili, 2014), which is a triangle set [^1] In this paper, we refer triangle set to a collection of connected or disconnected triangles. For those connected ones, they form a surface. with connected triangles formed by $k$ nearest neighbors ($k$-NNs). While triangle sets are effective in capturing location and orientation information, they often fall short in incorporating curvature knowledge, which is essential for accurate point cloud recognition (Sun et al., 2016; Czerniawski et al., 2016). Moreover, as depicted in the supplementary material, in certain $k$-NNs, the $k$ points may come from different surfaces of the object. These “noisy points” can lead to the distortion of $k$-NN triangle sets, significantly impacting the quality of the obtained geometric features (Ran et al., 2022).
@@
Moreover, our OPFR is backbone-agnostic, making it compatible with different 3D point clouds analysis architectures. We demonstrate its model-agnostic nature by adapting two representative backbones: PointNet++ (Qi et al., 2017b) and Point Transformer (Zhao et al., 2021b). It serves as an efficient plug-and-play module, and achieves substantial performance improvements. Empirical results prove its compatibility with different backbones. When incorporating with Point Transformer backbone, our OPFR achieves state-of-the-art performance for both point cloud classification and semantic segmentation tasks.
@@ 2.1. Deep Learning on Point Clouds
As PointNet++ establishes the hierarchical point clouds analysis framework, the focus of many works has shifted towards the development of local feature extractors, including convolution-based (Li et al., 2018; Jiang et al., 2018; Mao et al., 2019; Komarichev et al., 2019; Wu et al., 2019), attention-based (Guo et al., 2021; Zhao et al., 2021b; Yu et al., 2022), and graph-based (Yang et al., 2018; Te et al., 2018; Wang et al., 2019; Zhang et al., 2019; Xu et al., 2020) approaches. PointCNN (Li et al., 2018) learns a $\chi$-transformation from input point clouds, which attempts to re-organize inputs into canonical order. Subsequently, it utilizes vanilla convolution operations to extract local features. Point Transformer (Zhao et al., 2021b) replaces conventional shared-MLP modules with Transformer (Vaswani et al., 2017) blocks, serving as feature extractors within localized patch processing. DGCNN (Wang et al., 2019) utilizes dynamic graph structures to enhance feature learning and capture relationships between points. However, these works rely heavily on the learnability of feature extractors, potentially missing inherent local shape information. More recently, RepSurf (Ran et al., 2022) leverages triangle sets with connected triangles formed by $k$ nearest neighbors ($k$-NNs), to learn location and orientation-aware representations from geometric features explicitly. Although location and orientation features are explicitly injected into network architecture in RepSurf, the usage of curvature information still remains underexplored. Moreover, RepSurf relies on naive $k$-NNs to produce triangle sets and obtain geometric features, which are vulnerable to noisy points (Ran et al., 2022).
@@ 3. Methodology
The pipeline of On-the-fly Point Feature Representation (OPFR) is depicted in Fig. 1, and we illustrate the OPFR generating process for the right corner of table (highlighted in pink), which is one of our interested points. Firstly, we propose Hierarchical Sampling module, which takes each point in the point cloud as input and outputs several clusters (highlighted in blue) and corresponding centroids (highlighted in orange). This hierarchical sampling strategy improves the quality of triangle set for each point, thereby facilitating the development of subsequent geometric features. Then, for each point pair within the clusters, we design Curve Feature Generator module to generate geometric features, including location, orientation, and curvature. The inclusion of explicit curvature information allows us to more effectively capture the local geometry surrounding these point pairs. To enhance the efficiency and enable on-the-fly processing, we present Local Reference Constructor module. It approximates a local coordinate system (highlighted in red) for each point pair using adjacent points from triangle sets. Lastly, these obtained geometric features are further fed into shared-MLP followed by a pooling operation, constituting our final OPFR representation $\mathbf{r}_{i}$. The resultant OPFR representation $\mathbf{r}_{i}$ along with coordinate $\mathbf{x}_{i}$ can be directed into various point clouds analysis backbones, e.g., PointNet++ (Qi et al., 2017b) and Point Transformer (Zhao et al., 2021b), for end-to-end training.
@@ 3.3. Curve Feature Generator
PointTrans. (Zhao et al., 2021b) 1k pnts 93.7 90.6 82.3 80.7 5.187M 0.29G
@@ 3.4. On-the-fly Point Feature Representation
In Fig. 3, the three-view drawing depicts the OPFR values of 1-st channel for an airplane. The blue hues represent areas with smaller OPFR values, typically the airplane body, while the red hues indicate larger OPFR values, primarily associated with the airplane wings. This color differentiation underlines our OPFR is sensitive to curvature variation across the airplane’s structure, demonstrating the curvature-aware property of OPFR. We provide more visualization examples in the supplementary material. Additionally, it is important to highlight that, as shown in Fig. 4, our OPFR can outperform vanilla PFH by a large margin with the help of learnable shared-MLP. Furthermore, the introducing of shared-MLP only increases 0.012M learnable parameters, which is approximately negligible for most popular backbones (Qi et al., 2017b; Zhao et al., 2021b).
@@ 4. Experiments
We evaluate our OPFR on two primary tasks: point cloud classification and semantic segmentation. We choose two representative point cloud understanding models, PointNet++ (Qi et al., 2017b) and Point Transformer (Zhao et al., 2021b), as our backbones to evaluate the effectiveness and compatibility of OPFR representations across different backbone architectures. Additionally, we carry out ablation studies to demonstrate the effectiveness of our OPFR network designs and quantitatively evaluate the efficiency and quality of OPFR feature representations. Moreover, due to space constraints, we present qualitative results in the supplementary material.
@@ 4.1. Classification
Classification on ModelNet40. ModelNet40 (Wu et al., 2015) is one synthetic object classification benchmark, which contains 9843 training samples and 2468 testing samples. They contain 100 unique CAD models from 40 object categories. The experimental results are presented in Tab. 1. The results reveal that our OPFR significantly improves PointNet++ (Qi et al., 2017b) backbone by 3.8% OA and 3.2% mAcc, with just an additional 0.012M more parameters and 0.15G more FLOPs. The scaled-up OPFR further attains a slight improvement of 0.1% OA and 0.2% mAcc. Moreover, when integrated with transformer-based backbone, Point Transformer (Zhao et al., 2021b), our OPFR achieves the state-of-the-art 94.8% OA and 92.0% mAcc (+1.1% OA and +1.4% mAcc).
@@ 4.2. Semantic Segmentation
Semantic Segmentation on S3DIS. S3DIS (Armeni et al., 2016) encompasses 271 scenes which are distributed across 6 indoor areas, with each individual point being classified into one of 13 semantic labels. Following a common protocol (Tchapmi et al., 2017; Qi et al., 2017b), we evaluate the presented approach in two modes: (a) Area-5 is withheld for training and is used for testing, and (b) 6-fold cross-validation. In Tab. 2, our proposed OPFR considerably enhances PointNet++ (Qi et al., 2017b) by 14.7%/16.9%/3.0% (mIoU/mAcc/OA) on S3DIS 6-fold benchmark. Our result is comparable to PointNeXt-XL (Qian et al., 2022), with around 40$\times$ fewer parameters and 11$\times$ fewer FLOPs. When integrated with Point Transformer (Zhao et al., 2021b), the performance of OPFR exceeds previous state-of-the-art Superpoint Transformer (Sun et al., 2023) by 0.9%/0.1%/1.6% (mIoU/mAcc/OA) for S3DIS 6-fold. Meanwhile, on S3DIS Area-5, our OPFR attains mIoU/mAcc/OA of 72.6%/78.6%/91.7% (+2.2%/+2.1%/+0.9%), surpassing the prior state-of-the-art ConDaFormer (Duan et al., 2024).
@@ 5. Conclusion
We propose the novel plug-and-play module On-the-fly Point Feature Representation (OPFR) for various backbones. It explicitly captures local geometry including location, orientation and curvature through Curve Feature Generator module. We further develop the Local Reference Constructor module to improve efficiency and enable on-the-fly processing. Additionally, we introduce the Hierarchical Sampling module to mitigate the distortion of triangle sets that occurs in the naive $k$ nearest neighbors sampling, thereby enhancing the robustness of obtained geometric features. We evaluate the proposed OPFR on ModelNet40 and ScanObjectNN benchmarks for point cloud classification task, S3DIS for semantic segmentation task. For both PointNet++ and Point Transformer backbones, our presented OPFR achieves the state-of-the-art results on different benchmarks. The comprehensive empirical results demonstrate the backbone-agnostic nature of our proposed method. We believe that our work can prompt consideration of how to better leverage geometric knowledge in network architecture designs for understanding point clouds.